Can someone help me with this question? Given the pair of linear equations in two variables: Find the x- and y-intercepts (if any) for each line. Show your work. Apply elimination or substitution to find the coordinates of the point of intersection. State if there are no solutions or infinite solutions
x+5y=2
2x=-10y+4

Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

x+5y=2
2x=-10y+4
----------
The 2 equations are the same:
2x=-10y+4
2x + 10y = 4
x + 5y = 2
--> dependent, an infinite # of solutions.
--------------------
To find the intercepts:
for x, y = 0 --> x = 2
x-intercept at (2,0)
-------------
for y, x = 0 --> y = 2/5
y-intercept at (0,2/5)

Answer by checkley77(12844)

x+5y=2 multiply by 2 & subtract.
2x+10y=4
-(2x+10y=4)
-------------
0x+0y=0 this indicates the 2 lines are identical, thus there are infinite solutions.